OpenStudy (studygurl14):

check answer @kittiwitti1

OpenStudy (studygurl14):

15

OpenStudy (kittiwitti1):

I got a really weird mean... 6.625 o-o

OpenStudy (studygurl14):

lol

alones (alones):

LOL

OpenStudy (studygurl14):

You sure you did it right? @kittiwitti1

OpenStudy (kittiwitti1):

\[\frac{3+4+6+7+9+9+15}{8}=6.625\]

alones (alones):

Hrmm did you divide by 7?

OpenStudy (kittiwitti1):

... I left out the 11. fml

alones (alones):

to get mean

OpenStudy (studygurl14):

lolz

alones (alones):

Oh lolol

OpenStudy (kittiwitti1):

Also wot Doesn't adding a # make it 8 values

OpenStudy (studygurl14):

so...consensus...i'm right?

OpenStudy (kittiwitti1):

Are they saying new median or current/old median?

OpenStudy (studygurl14):

lol, if you added a number, the median would change to 8

OpenStudy (studygurl14):

3 + 4 + 6 + 7 + 9 + 9 + 11 + 15 = 64 64/8 = 8

OpenStudy (studygurl14):

thanks anyway, lol

OpenStudy (kittiwitti1):

I miscalculated the mean but yes, the new mean would be 8 and the new median would also be 8. You are correct;

OpenStudy (mathmate):

You can also start with the mean and median of the existing numbers, sum=49, mean=7 median=7, So what would you add to 3,4,6,7, ,9,9,11 ?

OpenStudy (mathmate):

@StudyGurl14 BTW, yes, 15 works.

OpenStudy (studygurl14):

lol thanks so much everyone!